Two small spheres each suspended by a thread of small length, L from a common support have masses m and 2m respectively. Each sphere carries a net charge of +Q. Assuming that the angle the thread make with the vertical is small, and the distance, X between the spheres can be given by the equation X = (5KQ^2. L / 2mg)^1/3
Where g = acceleration due to gravity, K = 1 / 40 and 0 = permittivity of free space. Show that the equation is homogeneous.
1
Expert's answer
2020-10-02T07:24:45-0400

Given equation is "X = (\frac{5KQ^2L}{2mg})^{1/3}"

where "K = \frac{1}{4\pi \epsilon_0}"


Equation will be homogeneous only when dimensions on both sides are same.

Since X is separation, so it dimension will be "[L]"

Dimension of K is "[ML^3T^{-4}A^{-2} ]"

Dimension of Q is "[AT]"

Dimension of L is [L]

Dimension of m is [M]

Dimension of g is "[LT^{-2}]"


Dimension of left hand side, "(\frac{[ML^3T^{-4}A^{-2}][A^2T^2][L] }{[M][LT^{-2}]})^{1/3} = [L]"


Since Dimension on both sides are same, so equation is homogeneous.